Suppose that X is a random variable from a binomial distribution with parameters n=12 and p. Consider the point estimate p̂=X/14

1. what's the bias of this estimate?
2. what is the value of the mean square error of this estimate if the actual value of p is 0.735
Answer #1

Solution: let Xro Bironinial (n=la, P) CCX) opa 12p varience (X) = $ np Ci-p) = 12x P(1-P) (a) ô = x/14 ulos P E Cx) = Elé):